how can i complete the square with:
x^2-10x=11
my answer, (-10x/2)^2 =11 was wrong.

Answer by nerdybill(7384)

x^2-10x=11
take 1/2 'b', square it and add to both sides:
x^2-10x+25=11+25
x^2-10x+25=36
(x-5)(x-5) = 36
(x-5)^2 = 36
x-5 = +-6
x = +-6 + 5
x = {-1, 11}